# A Robust Ensemble Algorithm for Ischemic Stroke Lesion Segmentation: Generalizability and Clinical Utility Beyond the ISLES Challenge

### v1

@@ Discussion
Choosing appropriate evaluation metrics can be difficult in challenge design. ISLES’22 addressed this by incorporating expert recommendations [maier2022metrics, reinke2023understanding]. As suggested, we employed a balanced mix of technical metrics (e.g., Dice scores) and clinically relevant metrics (e.g., number and volume of predicted ischemic lesions). This comprehensive approach allows for a broader assessment of solution performance and its readiness for real-world clinical applications, such as AI-assisted radiological reading.
The ISLES’22 challenge yielded fascinating insights into the landscape of stroke segmentation algorithms. Interestingly, the challenge leaderboard revealed that even algorithms based on similar CNN architectures and optimization strategies can exhibit variable performance. This reinforces the notion that factors beyond architecture, like hyper-parameter tuning, stochastic optimization, and training data sub-splitting (as in cross-validation), all contribute to model variability, even with a consistent dataset like ISLES’22 [haibe2020transparency]. However, the challenge also showcased the effectiveness of diverse algorithmic approaches. While achieving similar performance on most metrics, the top three ranked solutions employed different methodologies. The leading two teams utilized distinct CNN architectures (nnUnet [isensee2021nnu] and SegResNet[myronenko20193d]) and loss functions (Dice with binary cross-entropy vs. Dice with focal loss). Notably, the third-ranked solution adopted a completely different approach based on non-negative matrix factorization [ashtari2023factorizer]. This solution also leveraged the FLAIR modality (discarded by the top two teams), necessitating additional FLAIR-to-DWI co-registration. This remarkable diversity in successful solutions highlights the power of challenges in fostering innovation and the potential of ensemble methods that combine these strengths.
@@
Our ensemble algorithm demonstrates robust performance in handling a wide range of image and disease variations. This is evident from the successful generalization to unseen data from a new center, which achieved results similar to those of the trained center. Interestingly, while centers ##\##1 and ##\##3 (seen and unseen, respectively) showed similar metric distributions, a significant difference in Dice scores arose between centers ##\##1 and ##\##2 (both seen during training). This discrepancy can be attributed to two key factors. First, center ##\##2’s scans had considerably smaller stroke lesions. Second, these scans were acquired during the acute stroke phase. This observation highlights a crucial novel finding: the timing of brain imaging relative to stroke onset significantly impacts deep learning model performance. Lesion volumes from acute stroke scans were slightly overestimated compared to sub-acute ones. Despite this difference in Dice scores, the volumetric agreement between predictions and ground truth remained exceptionally high for both groups (Pearson’s r = 0.99 and 0.98 for acute and sub-acute, respectively). Ischemia detectability, as measured by lesion-wise F1 scores, also remained statistically similar between groups. These results lead to two key conclusions: $i)$ The model demonstrates remarkable generalizability across different stroke phases (acute and sub-acute) and imaging centers. This suggests the successful capture of stroke lesion variability, avoiding reliance on center-specific MRI features. $ii)$ The stroke phase at scan acquisition demonstrably influences the model. This is understandable as early (acute) scans exhibit different MR characteristics compared to later ones (sub-acute) due to evolving tissue changes. This aligns with established knowledge about how DWI and ADC values fluctuate with stroke progression [fung2011mr]. Similarly, DWI sensitivity (specificity) ranges between 73% (92%) 3 hours from the stroke event to 92% (97%) 12 hours from the stroke event [chalela2007magnetic, merino2010imaging]. Furthermore, false negatives may also increase with early DWI acquisition [sylaja2008expect, merino2010imaging].
@@
Beyond lesion volume, our model sheds light on stroke etiology. Traditionally, stroke type and affected vascular territories are crucial for determining the underlying cause, impacting treatment decisions and prevention strategies (e.g., Merino et al.[merino2010imaging], Kim et al.[kim2014magnetic]). Existing research establishes associations between DWI lesion patterns and stroke causes (e.g. Kang et al.[kang2003association]). For instance, scattered infarcts across multiple territories, single cortico-subcortical lesions, or multiple lesions in the anterior and posterior circulation often indicate cardioembolism [cho2007mechanism, kim2014magnetic, merino2010imaging, kang2003association]. Conversely, large artery atherosclerosis typically presents with lesions in a single vascular territory [merino2010imaging, kang2003association]. In this context, our deep learning algorithm stands out by accurately characterizing DWI images from a multi-faceted perspective. The model effectively segments ischemia in different stroke sub-groups (SVI, scattered infarcts, and mixed SVI/scattered) with consistent performance. Importantly, the model tackles even small ischemic volumes (e.g., embolic showers) with high detectability (lesion-wise F1 scores comparable to larger lesions). Furthermore, it accurately identifies stroke subgroups and affected vascular territories using predicted lesion masks (multi-class balanced accuracies of $\sim$ 87% and $\sim$ 98%, respectively, Table 2). These findings suggest that the algorithm’s outputs extend beyond lesion volume, providing valuable clinical insights into stroke type and underlying cause, ultimately aiding downstream clinical assessments.
To further assess segmentation quality, we employed a blinded evaluation similar to a Turing test. Nine experienced neuroradiologists from various institutions judged the completeness (lesion identification) and correctness (delineation accuracy) of both manual and algorithmic segmentations. Their preference, aligning with Kofler et al.’s findings for brain tumors [kofler2021we], consistently favored the ensemble algorithm’s outputs. This finding suggests that deep learning algorithms can match or even surpass human experts in identifying brain infarcts on MRI scans.
@@ Methods
In order to devise an automatic algorithm that can identify brain ischemia under heterogeneous real-world imaging scenarios, the public challenge ISLES’22 is organized. The ISLES’22 challenge enables a fast and extended benchmarking of algorithmic strategies for tackling the task. Therefore, this section is organized in two parts. The first section explains how the challenge is structured. The BIAS (Biomedical Image Analysis ChallengeS) [maier2020bias] guideline is followed. The second section details how the identified algorithms are evaluated to come up with a clinically useful solution.
@@
The aims, structure, and organization of the challenge are available for the teams several weeks before the dataset is released. A detailed description of the challenge organization is available in de la Rosa et al.[islesdocument]. The challenge is organized in three phases: a train, a sanity-check, and a test phase. In the train phase, participating teams have six weeks to develop a solution to the task using a labeled MRI dataset. All teams have access to the data at the same time. There are no technical constraints on the employed algorithmic solution. In the sanity-check phase, participants can test their devised Docker solutions over a few train set scans. The aim of this phase is to guarantee that each team’s algorithm properly works in the challenge organizers’ servers. Therefore, multiple submissions of the algorithms are allowed. In the test phase, participating teams are requested to submit a Docker containing their final algorithmic solution. Teams can submit just a single time to this phase. The algorithm is later run by the challenge organizers over the hidden test set. Algorithmic performance is measured by computing relevant metrics (below detailed) using the predicted segmentations and the ground truth masks. Later, teams are ranked based on their yield performance metrics.
@@
The dataset used in this study is customarily devised for the purposes of the challenge. It contains multi-center and multi-scanner data, and it consists of MRI scans (n=400) acquired during the early/late acute and the early sub-acute stroke phase of patients across three European health centers. The train (n=250) and test (n=150) sets include scans from two and three healthcare centers, respectively. As the challenge was focused on post-operative stroke images, the inclusion criteria consisted on patients who underwent MRI imaging after receiving mechanical thrombectomy treatment. Furthermore, as a proof of concept, we aim to test the generalization capability of the devised algorithms over a small, single-center subset of hyper-acute/early acute stroke scans (12.5%, n=50) before intervention, which represents a third part of the test set. Table 4 provides a summary of the ISLES’22 dataset. All cases include DWI (b-value=1000 s/mm2), FLAIR, and ADC MRI. The ischemic stroke segmentation ground truths are obtained using an algorithm-human hybrid iterative method and are later revised and refined by one out of three experienced neuroradiologists with more than 10 years of experience (RW, JSK and BW). The MRI images are released in their native acquisition space (ground truth masks are released in the DWI/ADC space) after minimal pre-processing. Thus, pre-processing is solely performed with the purpose of patient de-identification and therefore consists of MRI skull-stripping. The reason for releasing the dataset ‘as raw as possible’ is to encourage the development of algorithms that could deal with real-world raw imaging data, which suffers from a large variability (signal-to-noise, resolutions, variable parameter settings, etc.) and therefore has its own technical limitations (e.g., different MR modalities, as FLAIR and DWI, are not co-registered). In this sense, participants are also challenged to devise end-to-end algorithms that can deal with the pre-processing and curation of the images. For further details about this dataset please check the corresponding data descriptor [hernandez2022isles].
@@
ISLES’22[hernandez2022isles] Train 250 200 / 50 / 0 MT 0 250 Sub-acute
@@
JHCSC[liu2023large] Test 1686 - ivtPA 809 870 Acute and sub-acute
Table 4: Data summary. The ISLES’22 dataset was used in the challenge competition, while the JHCSC was used as an external testing dataset in a post-challenge setting. JHCSC: Johns Hopkins Comprehensive Stroke Center; RT: reperfusion treatment; MT: mechanical thrombectomy; ivtPA: intravenous tissue plasminogen activator. Further details about the datasets are available in the corresponding data descriptors [hernandez2022isles, liu2023large].
@@
The algorithmic results are evaluated by comparing the predicted lesions masks with the (manually traced) ground truth masks. Selection of the appropriate metrics to assess the method’s performance was heavily discussed within the challenge organization team based on current recommendation guidelines [maier2022metrics, reinke2023understanding]. The aim was to conduct a clinically relevant evaluation unbiased by diverse lesion characteristics and lesion patterns. The following metrics that are well known in the literature for both the research and radiological communities were included: Dice score [dice1945measures] (DSC), the absolute volume difference ($\mathrm{AVD=|Volume_{\,predicted}-Volume_{\,groundtruth}|}$), lesion-wise F1 score (defined as in Section Statistical analysis by considering instance lesions), and the absolute lesion count difference ($\mathrm{ALD=|\#\,Lesions_{\,predicted}-\#\,Lesions_{\,ground\>truth}|}$). Lesion-wise metrics were computed after isolating disconnected ischemias in the binary masks through connected-component analysis. Note that including complementary segmentation and detection metrics is beneficial and helps evaluate models in a robust, subject, and lesion-wise unbiased scheme [sudre2022valdo]. Dice was used to evaluate semantic segmentation at the entire subject level. Given that segmentation metrics (as DSC) do not provide insights about the clinical utility of the predictions, we also included a general clinical metric of interest as the AVD. Furthermore, lesion detection metrics (which are computed for each instance lesion) were considered since the dataset has scans with multiple and small emboli, which might not drive important changes in the DSC and AVD metrics under the presence of large disconnected infarcts. For example, algorithms missing a small lesion can achieve a very high DSC if a much larger infarct within the same scan is properly segmented. However, an algorithm detecting both lesions (which is clinically desired) might still perform worse in terms of DSC and AVD. Consequently, to penalize algorithms that under- or over-estimate the number of predicted lesions, we evaluate the algorithm’s capability to detect each (single) disconnected ischemia through the lesion-wise F1 score and ALD. Implementation details of the four metrics can be found in the challenge Python notebooks[islesrepo, atlasrepo].
@@
The final competition ranking is obtained from the test phase, which is blind (there is no access for the teams to the MRI images to be predicted) and single-shoot (one submission per team is solely allowed), thus completely precluding participants from any sort of overfitting strategy. As similarly done in previous ISLES editions[maier2017isles, winzeck2018isles, hakim2021predicting], the ranking is performed in a ‘rank then aggregate’ fashion [maier2018rankings]. In a nutshell, the ranking scheme is obtained as the average ranking position obtained from all the individual rankings by case and by metric. A thousand bootstraps are conducted by repeatedly drawing samples with replacements and recalculating the rankings in each iteration. Furthermore, as a complementary analysis, we calculated the challenge ranking through a thousand bootstraps ‘aggregate then rank’ scheme.
@@
In this experiment, we evaluate if the deep learning models can identify the affected brain vascular territory in the MRI scans. For doing so, the test-phase scans are linearly registered to a FLAIR MNI template with vascular territory annotations [schirmer2019spatial] using NiftyReg [ourselin2002robust]. Later, the lesion load over each vascular territory is quantified using the ground truth annotations and each scan receives a label of the vascular territory that yields the largest lesion load. The considered vascular territories are the ones covered by the middle cerebral artery, the anterior cerebral artery, the posterior cerebral artery, the arteries perfusing the cerebellum, and the ones perfusing the pons and medulla. The deep learning predictions of the vascular territories are generated by finding the vascular territory with the largest (predicted) lesion volume. Then, we assess through classification metrics how well the algorithms identify the stroke vascular territory.
@@
Nine radiologists from three healthcare centers (University Hospital of the Technical University Munich, University Medical Center Hamburg-Eppendorf, and University Hospital of Bern) blindly rated the quality of the ischemic stroke masks generated either by experts or by the ensemble algorithm. Forty or forty-one scans with three annotated slices each (two axial, one sagittal) were provided to each radiologist. All images were randomized, and no information about the annotator (human or algorithm) was provided. Radiologists were asked to rate the completeness of the segmented lesion and the correctness of their contours on a 1-6 scale as similarly done in Kofler et al.[kofler2021we] (see Supplementary material S6 for the criteria used).
@@
The ensemble algorithm is tested over a public, external, ischemic stroke cohort (n = 1686) [liu2023large] including raw MRI (such as DWI, ADC, FLAIR, etc.), patient (e.g., age, sex, race) and clinical data (e.g., treatment, NIHSS and mRS scores, etc.). Table 4 summarises the dataset characteristics. Images were acquired over ten years using eleven 1.5T or 3T scanners from the four major machine vendors (Siemens, GE, Toshiba, and Philips). NIHSS and mRS scores were respectively performed at patient admission and at 90-day follow-up. Moreover, the time from symptom onset to MRI acquisition was recorded when the patient or the caregiver was confident about symptom onset. MRI was mostly performed six or more hours from symptom onset, before or after administration of intravenous tissue plasminogen activator. To predict ischemic lesions with the ensemble algorithm, all scans were priorly skull-stripped using HD-BET [isensee2019automated].
@@
The ISLES’22 dataset used for the challenge is open and freely available under the Creative Commons CC BY 4.0 license. The train dataset is available in www.zenodo.org/[islesdata]. The external stroke dataset used for validating the ensemble algorithm is available through ICPSR [icpsr, liu2023large].
@@
Performance metrics are available in Table 1 and also through https://grand-challenge.org/[GC_isles]. Note that this challenge continues accepting submissions and, therefore, the online leaderboard is constantly getting updated. In this study, only the solutions received for the MICCAI’2022 challenge are evaluated.
@@
To help participants get familiar with the data and with the challenge performance metrics a Python notebook [islesrepo] was made available in advance. Moreover, in order to facilitate the teams during the algorithms’ submission process, a Docker template and a Docker creation tutorial were shared with the teams [dockerisles]. Figures of this work were generated in R [computing2013r] using the ggplot2[hadley2016ggplot2] and the patchwork[patchwork] software packages. Figure 5 was generated with 3D Slicer (https://www.slicer.org/) [fedorov20123d].

### v2

@@ Discussion
Choosing appropriate evaluation metrics can be difficult in challenge design. ISLES’22 addressed this by incorporating expert recommendations [38, 39]. As suggested, we employed a balanced mix of technical metrics (e.g., Dice scores) and clinically relevant metrics (e.g., number and volume of predicted ischemic lesions). This comprehensive approach allows for a broader assessment of solution performance and its readiness for real-world clinical applications, such as AI-assisted radiological reading.
The ISLES’22 challenge yielded fascinating insights into the landscape of stroke segmentation algorithms. Interestingly, the challenge leaderboard revealed that even algorithms based on similar CNN architectures and optimization strategies can exhibit variable performance. This reinforces the notion that factors beyond architecture, like hyper-parameter tuning, stochastic optimization, and training data sub-splitting (as in cross-validation), all contribute to model variability, even with a consistent dataset like ISLES’22 [42]. However, the challenge also showcased the effectiveness of diverse algorithmic approaches. While achieving similar performance on most metrics, the top three ranked solutions employed different methodologies. The leading two teams utilized distinct CNN architectures (nnUnet [26] and SegResNet[30]) and loss functions (Dice with binary cross-entropy vs. Dice with focal loss). Notably, the third-ranked solution adopted a completely different approach based on non-negative matrix factorization [32]. This solution also leveraged the FLAIR modality (discarded by the top two teams), necessitating additional FLAIR-to-DWI co-registration. This remarkable diversity in successful solutions highlights the power of challenges in fostering innovation and the potential of ensemble methods that combine these strengths.
@@
Our ensemble algorithm demonstrates robust performance in handling a wide range of image and disease variations. This is evident from the successful generalization to unseen data from a new center, which achieved results similar to those of the trained center. Interestingly, while centers ##\##1 and ##\##3 (seen and unseen, respectively) showed similar metric distributions, a significant difference in Dice scores arose between centers ##\##1 and ##\##2 (both seen during training). This discrepancy can be attributed to two key factors. First, center ##\##2’s scans had considerably smaller stroke lesions. Second, these scans were acquired during the acute stroke phase. This observation highlights a crucial novel finding: the timing of brain imaging relative to stroke onset significantly impacts deep learning model performance. Lesion volumes from acute stroke scans were slightly overestimated compared to sub-acute ones. Despite this difference in Dice scores, the volumetric agreement between predictions and ground truth remained exceptionally high for both groups (Pearson’s r = 0.99 and 0.98 for acute and sub-acute, respectively). Ischemia detectability, as measured by lesion-wise F1 scores, also remained statistically similar between groups. These results lead to two key conclusions: $i)$ The model demonstrates remarkable generalizability across different stroke phases (acute and sub-acute) and imaging centers. This suggests the successful capture of stroke lesion variability, avoiding reliance on center-specific MRI features. $ii)$ The stroke phase at scan acquisition demonstrably influences the model. This is understandable as early (acute) scans exhibit different MR characteristics compared to later ones (sub-acute) due to evolving tissue changes. This aligns with established knowledge about how DWI and ADC values fluctuate with stroke progression [12]. Similarly, DWI sensitivity (specificity) ranges between 73% (92%) 3 hours from the stroke event to 92% (97%) 12 hours from the stroke event [11, 4]. Furthermore, false negatives may also increase with early DWI acquisition [43, 4].
@@
Beyond lesion volume, our model sheds light on stroke etiology. Traditionally, stroke type and affected vascular territories are crucial for determining the underlying cause, impacting treatment decisions and prevention strategies (e.g., Merino et al.[4], Kim et al.[44]). Existing research establishes associations between DWI lesion patterns and stroke causes (e.g. Kang et al.[13]). For instance, scattered infarcts across multiple territories, single cortico-subcortical lesions, or multiple lesions in the anterior and posterior circulation often indicate cardioembolism [45, 44, 4, 13]. Conversely, large artery atherosclerosis typically presents with lesions in a single vascular territory [4, 13]. In this context, our deep learning algorithm stands out by accurately characterizing DWI images from a multi-faceted perspective. The model effectively segments ischemia in different stroke sub-groups (SVI, scattered infarcts, and mixed SVI/scattered) with consistent performance. Importantly, the model tackles even small ischemic volumes (e.g., embolic showers) with high detectability (lesion-wise F1 scores comparable to larger lesions). Furthermore, it accurately identifies stroke subgroups and affected vascular territories using predicted lesion masks (multi-class balanced accuracies of $\sim$ 87% and $\sim$ 98%, respectively, Table 2). These findings suggest that the algorithm’s outputs extend beyond lesion volume, providing valuable clinical insights into stroke type and underlying cause, ultimately aiding downstream clinical assessments.
To further assess segmentation quality, we employed a blinded evaluation similar to a Turing test. Nine experienced neuroradiologists from various institutions judged the completeness (lesion identification) and correctness (delineation accuracy) of both manual and algorithmic segmentations. Their preference, aligning with Kofler et al.’s findings for brain tumors [40], consistently favored the ensemble algorithm’s outputs. This finding suggests that deep learning algorithms can match or even surpass human experts in identifying brain infarcts on MRI scans.
@@ Methods
In order to devise an automatic algorithm that can identify brain ischemia under heterogeneous real-world imaging scenarios, the public challenge ISLES’22 is organized. The ISLES’22 challenge enables a fast and extended benchmarking of algorithmic strategies for tackling the task. Therefore, this section is organized in two parts. The first section explains how the challenge is structured. The BIAS (Biomedical Image Analysis ChallengeS) [23] guideline is followed. The second section details how the identified algorithms are evaluated to come up with a clinically useful solution.
@@
The aims, structure, and organization of the challenge are available for the teams several weeks before the dataset is released. A detailed description of the challenge organization is available in de la Rosa et al.[25]. The challenge is organized in three phases: a train, a sanity-check, and a test phase. In the train phase, participating teams have six weeks to develop a solution to the task using a labeled MRI dataset. All teams have access to the data at the same time. There are no technical constraints on the employed algorithmic solution. In the sanity-check phase, participants can test their devised Docker solutions over a few train set scans. The aim of this phase is to guarantee that each team’s algorithm properly works in the challenge organizers’ servers. Therefore, multiple submissions of the algorithms are allowed. In the test phase, participating teams are requested to submit a Docker containing their final algorithmic solution. Teams can submit just a single time to this phase. The algorithm is later run by the challenge organizers over the hidden test set. Algorithmic performance is measured by computing relevant metrics (below detailed) using the predicted segmentations and the ground truth masks. Later, teams are ranked based on their yield performance metrics.
@@
The dataset used in this study is customarily devised for the purposes of the challenge. It contains multi-center and multi-scanner data, and it consists of MRI scans (n=400) acquired during the early/late acute and the early sub-acute stroke phase of patients across three European health centers. The train (n=250) and test (n=150) sets include scans from two and three healthcare centers, respectively. As the challenge was focused on post-operative stroke images, the inclusion criteria consisted on patients who underwent MRI imaging after receiving mechanical thrombectomy treatment. Furthermore, as a proof of concept, we aim to test the generalization capability of the devised algorithms over a small, single-center subset of hyper-acute/early acute stroke scans (12.5%, n=50) before intervention, which represents a third part of the test set. Table 4 provides a summary of the ISLES’22 dataset. All cases include DWI (b-value=1000 s/mm2), FLAIR, and ADC MRI. The ischemic stroke segmentation ground truths are obtained using an algorithm-human hybrid iterative method and are later revised and refined by one out of three experienced neuroradiologists with more than 10 years of experience (RW, JSK and BW). The MRI images are released in their native acquisition space (ground truth masks are released in the DWI/ADC space) after minimal pre-processing. Thus, pre-processing is solely performed with the purpose of patient de-identification and therefore consists of MRI skull-stripping. The reason for releasing the dataset ‘as raw as possible’ is to encourage the development of algorithms that could deal with real-world raw imaging data, which suffers from a large variability (signal-to-noise, resolutions, variable parameter settings, etc.) and therefore has its own technical limitations (e.g., different MR modalities, as FLAIR and DWI, are not co-registered). In this sense, participants are also challenged to devise end-to-end algorithms that can deal with the pre-processing and curation of the images. For further details about this dataset please check the corresponding data descriptor [37].
@@
ISLES’22[37] Train 250 200 / 50 / 0 MT 0 250 Sub-acute
@@
JHCSC[41] Test 1686 - ivtPA 809 870 Acute and sub-acute
Table 4: Data summary. The ISLES’22 dataset was used in the challenge competition, while the JHCSC was used as an external testing dataset in a post-challenge setting. JHCSC: Johns Hopkins Comprehensive Stroke Center; RT: reperfusion treatment; MT: mechanical thrombectomy; ivtPA: intravenous tissue plasminogen activator. Further details about the datasets are available in the corresponding data descriptors [37, 41].
@@
The algorithmic results are evaluated by comparing the predicted lesions masks with the (manually traced) ground truth masks. Selection of the appropriate metrics to assess the method’s performance was heavily discussed within the challenge organization team based on current recommendation guidelines [38, 39]. The aim was to conduct a clinically relevant evaluation unbiased by diverse lesion characteristics and lesion patterns. The following metrics that are well known in the literature for both the research and radiological communities were included: Dice score [46] (DSC), the absolute volume difference ($\mathrm{AVD=|Volume_{\,predicted}-Volume_{\,groundtruth}|}$), lesion-wise F1 score (defined as in Section Statistical analysis by considering instance lesions), and the absolute lesion count difference ($\mathrm{ALD=|\#\,Lesions_{\,predicted}-\#\,Lesions_{\,ground\>truth}|}$). Lesion-wise metrics were computed after isolating disconnected ischemias in the binary masks through connected-component analysis. Note that including complementary segmentation and detection metrics is beneficial and helps evaluate models in a robust, subject, and lesion-wise unbiased scheme [47]. Dice was used to evaluate semantic segmentation at the entire subject level. Given that segmentation metrics (as DSC) do not provide insights about the clinical utility of the predictions, we also included a general clinical metric of interest as the AVD. Furthermore, lesion detection metrics (which are computed for each instance lesion) were considered since the dataset has scans with multiple and small emboli, which might not drive important changes in the DSC and AVD metrics under the presence of large disconnected infarcts. For example, algorithms missing a small lesion can achieve a very high DSC if a much larger infarct within the same scan is properly segmented. However, an algorithm detecting both lesions (which is clinically desired) might still perform worse in terms of DSC and AVD. Consequently, to penalize algorithms that under- or over-estimate the number of predicted lesions, we evaluate the algorithm’s capability to detect each (single) disconnected ischemia through the lesion-wise F1 score and ALD. Implementation details of the four metrics can be found in the challenge Python notebooks[48, 49].
@@
The final competition ranking is obtained from the test phase, which is blind (there is no access for the teams to the MRI images to be predicted) and single-shoot (one submission per team is solely allowed), thus completely precluding participants from any sort of overfitting strategy. As similarly done in previous ISLES editions[21, 50, 51], the ranking is performed in a ‘rank then aggregate’ fashion [14]. In a nutshell, the ranking scheme is obtained as the average ranking position obtained from all the individual rankings by case and by metric. A thousand bootstraps are conducted by repeatedly drawing samples with replacements and recalculating the rankings in each iteration. Furthermore, as a complementary analysis, we calculated the challenge ranking through a thousand bootstraps ‘aggregate then rank’ scheme.
@@
In this experiment, we evaluate if the deep learning models can identify the affected brain vascular territory in the MRI scans. For doing so, the test-phase scans are linearly registered to a FLAIR MNI template with vascular territory annotations [52] using NiftyReg [53]. Later, the lesion load over each vascular territory is quantified using the ground truth annotations and each scan receives a label of the vascular territory that yields the largest lesion load. The considered vascular territories are the ones covered by the middle cerebral artery, the anterior cerebral artery, the posterior cerebral artery, the arteries perfusing the cerebellum, and the ones perfusing the pons and medulla. The deep learning predictions of the vascular territories are generated by finding the vascular territory with the largest (predicted) lesion volume. Then, we assess through classification metrics how well the algorithms identify the stroke vascular territory.
@@
Nine radiologists from three healthcare centers (University Hospital of the Technical University Munich, University Medical Center Hamburg-Eppendorf, and University Hospital of Bern) blindly rated the quality of the ischemic stroke masks generated either by experts or by the ensemble algorithm. Forty or forty-one scans with three annotated slices each (two axial, one sagittal) were provided to each radiologist. All images were randomized, and no information about the annotator (human or algorithm) was provided. Radiologists were asked to rate the completeness of the segmented lesion and the correctness of their contours on a 1-6 scale as similarly done in Kofler et al.[40] (see Supplementary material S6 for the criteria used).
@@
The ensemble algorithm is tested over a public, external, ischemic stroke cohort (n = 1686) [41] including raw MRI (such as DWI, ADC, FLAIR, etc.), patient (e.g., age, sex, race) and clinical data (e.g., treatment, NIHSS and mRS scores, etc.). Table 4 summarises the dataset characteristics. Images were acquired over ten years using eleven 1.5T or 3T scanners from the four major machine vendors (Siemens, GE, Toshiba, and Philips). NIHSS and mRS scores were respectively performed at patient admission and at 90-day follow-up. Moreover, the time from symptom onset to MRI acquisition was recorded when the patient or the caregiver was confident about symptom onset. MRI was mostly performed six or more hours from symptom onset, before or after administration of intravenous tissue plasminogen activator. To predict ischemic lesions with the ensemble algorithm, all scans were priorly skull-stripped using HD-BET [54].
@@
The ISLES’22 dataset used for the challenge is open and freely available under the Creative Commons CC BY 4.0 license. The train dataset is available in www.zenodo.org/[58]. The external stroke dataset used for validating the ensemble algorithm is available through ICPSR [59, 41].
@@
Performance metrics are available in Table 1 and also through https://grand-challenge.org/[24]. Note that this challenge continues accepting submissions and, therefore, the online leaderboard is constantly getting updated. In this study, only the solutions received for the MICCAI’2022 challenge are evaluated.
@@
To help participants get familiar with the data and with the challenge performance metrics a Python notebook [48] was made available in advance. Moreover, in order to facilitate the teams during the algorithms’ submission process, a Docker template and a Docker creation tutorial were shared with the teams [60]. Figures of this work were generated in R [61] using the ggplot2[62] and the patchwork[63] software packages. Figure 5 was generated with 3D Slicer (https://www.slicer.org/) [64].